# Measuring temporal effects of agent knowledge by date-controlled tool use

### v1

@@ Abstract
Temporal progression is an integral part of knowledge accumulation and update. Web search is frequently adopted as grounding for agent knowledge, yet its inappropriate configuration affects the quality of agent responses. Here, we construct a tool-based out-of-sample testing framework to measure the knowledge variability of large language model (LLM) agents from distinct date-controlled tools (DCTs). We demonstrate the temporal effects of an LLM agent as a writing assistant, which can use web search to help complete scientific publication abstracts. We show that temporal effects of the search engine translates into tool-dependent agent performance but can be alleviated with base model choice and explicit reasoning instructions such as chain-of-thought prompting. Our results indicate that agent evaluation should take a dynamical view and account for the temporal influence of tools and the updates of external resources [^1] The code and datasets used for the work are available at https://github.com/RealPolitiX/agent_oost..
@@ 1 Introduction
AI agents based on LLMs and equipped with tools (Mialon et al., 2023; Wang et al., 2024) are well-suited for complex real-world tasks (Gao et al., 2024; Xu et al., 2024) because of their extended capabilities. Their potential to become virtual assistants, paraprofessionals, or “copilots" holds promise for improving the productivity and creativity of the scientific, medical workforces and beyond (Wachter and Brynjolfsson, 2024; Wornow et al., 2024; Bousetouane, 2025). The evaluation standards for AI agents are still in flux (Kapoor et al., 2024; Højmark et al., 2024) and they are urgently needed in specialized domains and realistic scenarios where the the outcomes convey greater bearing on their adoption. Recent works demonstrated the feasibility of LLMs in predicting temporal events (Ye et al., 2024a) and carry out time series forecasting (Tang et al., 2024), but their equivalents in agentic systems are not yet realized. Scientific knowledge and claims have a strong temporal dependence but they have so far been less studied in the context of generative language models (Zhao et al., 2024; Park et al., 2024). We devised a proxy task intended for measuring the agent’s usability as a writing assistant with access to external sources and cast the problem as abstract completion for scientific publications (see Fig. 1a).
@@
Web search is an essential tool for grounding agent knowledge to the current and bygone worlds (Pavlick, 2023) and appears in many applications as capability extension (Zhou et al., 2024a; Song et al., 2024). Nevertheless, it is subject to the recency and primacy bias of the search engine (Lawrence and Giles, 1998) and the cognitive bias of the users who collect information (Lau and Coiera, 2007). The term search engine manipulation effect (Epstein and Robertson, 2015) was coined to refer to the search results’ influence on public opinions of societal issues. Independent of search engines, factual and scientific knowledge also experience constant updates over time (Arbesman, 2013).
@@
Figure 1: (a) Illustration of the OOS testing framework for agent knowledge, $t_{p}$ indicates the time of publication (b) Temporal tool selection in a ReAct-style agent that performs text completion task in (a) with a selected tool.
@@
While temporal generalization remains challenging for language models (Lazaridou et al., 2021; Wallat et al., 2024), explicit tuning of time-related tool parameters in LLM agents can offer an alternative way to reduce temporal effects of the base model. These effects are also potentially sources of robustness and reproducibility issues of the agentic system that warrant investigation (Ye et al., 2024b). Date control in reality can manifest passively because the tool-interfaced computer programs have an intrinsic time stamp or a versioned release over time (Zhang et al., 2009). Alternatively, date control can be imposed actively because of copyright, paywall, or local policy. Content access in the past can be controlled retroactively as policy changes (Aral and Dhillon, 2021).
@@
Out-of-sample (OOS) testing is the ultimate test for temporal prediction methods (Hansen and Timmermann, 2015). Analogous OOS assessments in the text domain include predicting future events (Ye et al., 2024a) or generating hypotheses (Zhou et al., 2024b) conditioned on past knowledge. We investigate the comparable problem from the tool use perspective, where the agent has access to changing internet-scale information. Because scientific breakthroughs often lead to significant knowledge updates, they are good markers for temporal knowledge progression [^2] Although the judgement on breakthroughs are ultimately subjective and can change with time, our motivation to use them is because of their noticeable footprints on the internet.. In this work, we aim to investigate the following research questions:
@@ 2. RQ2:
    Can the agent determine which date-controlled version of the tool to use for a task?
Our contributions are: (i) Formulation of tool-based OOS test for time-dependent knowledge for LLM agents; (ii) Introduction of the SciBreak dataset containing the publication records of public-endorsed scientific breakthroughs from 2000 to 2024. (iii) Investigation of the temporal effects of LLM agent performance and behavior. Besides, we also discuss the impact of temporal information on the agent capability and usability and its implications to subsequent research and applications.
@@ 2 Related works
Previous works have shown that the latent space of LLMs has a direction of time (Gurnee and Tegmark, 2024). Recent investigations show that model performance is affected by the lack of temporal grounding in the pre-training process (Zhao et al., 2024), which can hinder the elicitation of appropriate time-sensitive knowledge at task execution. LLMs have been used for predicting future events (Ye et al., 2024a) and conducting time series analysis (Tang et al., 2024; Jin et al., 2024). Previous works have shown LLMs often struggle with tasks that require a consistent temporal understanding (Qiu et al., 2024). This limitation can be improved with techniques such as temporally-informed chain-of-thought (CoT) prompting (Xiong et al., 2024).
@@
3 Tool-based OOS testing
@@
Definition 3.3 (Tool-based OOS test).
A performance test that induces OOS condition by adjusting the tool parameters of an agentic system. A temporal version of the test alters the time information of tools and therefore measures the reliability of agent performance under such conditions.
@@ 4 Testing framework implementation
We constructed a dataset, SciBreak, with a clear time-delimited footprint on the internet—scientific breakthroughs. The dataset is partly based on peer-reviewed publications collated and categorized in Wuestman et al. (2020), including the annual top-ten-ranked scientific breakthroughs from mid-1990s till 2012 collated by the journal Science (publications not limited to one journal) at the end of each year. We chose records from year 2000 to 2012 and extended to year 2024. Each year contains up to $\sim$ 20 publications, with multiple contributing to one breakthrough. We collected the abstracts of the associated publications through web scraping from the public databases PubMed [^3] https://pubmed.ncbi.nlm.nih.gov/ and SAO/NASA ADS Abstract Service [^4] https://ui.adsabs.harvard.edu/.
Agent construction and configuration
We combined DCT with the ReAct agent architecture (Yao et al., 2023) (ReAct+DCT) which allows interleaved thinking and action. The agents were constructed from closed-source models, including OpenAI’s GPT-3.5-turbo (gpt-3.5-turbo-0125), GPT-4-turbo (gpt-4-turbo-2024-04-09), and GPT-4o (gpt-4o-2024-08-06) as the base model.
@@
Table 1: Performance of ReAct-style LLM agents on text completion ($\gamma=0.5$) following web search with DCTs. For the row of Input, the metrics are computed between the input and the ground truth.
@@
5 Results
We analyzed the agent performance and reasoning patterns, following the previous taxonomy of LLM temporal behaviors in Wallat et al. (2024).
Reasoning about time
In our evaluation experiments, the agent reasoning behavior related to its awareness of time (e.g. does the tool have time-appropriate utility?) is triggered in two scenarios: (i) When the web search returns nothing or little relevant information to assist task completion. The agent then proceeds to complete the task with the internal knowledge of the base LLM. (ii) In temporal tool selection, when the agent is given an explicit CoT prompt to direct its reasoning towards considering the relavance of information to the topic.
Evaluation results in Table 2 show that the STS increases by including CoT prompting (ReAct + CoT) than with ReAct only, when the agent by default selects $\mathcal{T}_{t_{+}-1,t_{+}}$ as the tool. Here, $t\in[t_{-},t_{+}]$ being the date range of the tools. The agent then explores more date choices driven by its internal understanding of the scientific concepts present in the input paragraph. These behavioral characteristics allow the agent to handle non-existent and underspecified contexts in the OOS setting. The performance boost of ReAct + CoT agent pattern requires a model with sufficient reasoning capability such as GPT-4, while for GPT-3.5, it is more prone to failure and the performance gain is reversed.
Agent model Agent pattern Pub. in 2003 Pub. in 2015
Rouge-L STS Rouge-L STS
GPT-3.5 -turbo ReAct 0.447 0.732 0.447 0.794
ReAct + CoT 0.487 0.640 0.518 0.607
GPT-4 -turbo ReAct 0.635 0.888 0.588 0.783
ReAct + CoT 0.649 0.897 0.644 0.859
Temporal effects across models and masking
High-capacity models with pronounced reasoning capabilities are capable of examining tool dates. The task evaluated the model capability with increasing level of word masking. In Tables 1-4, the outcome contains two major trends: (i) The more advanced models can recover more of the missing semantic content in the masked input, as indicated by the significant increase of STS from LLM agents based on GPT-3.5 to GPT-4o. (ii) There is noticeable variability of agent performance between knowledge generated more recently than before 2010. (iii) For the same model, varying the masking ratio of input largely preserves the date sensitivity in the performance. Similar time-dependent performance change has been described in a different context for LLMs (Zhao et al., 2024). Overall, the temporal effects are less severe in more capable models than their lesser counterparts.
6 Discussion
Agent vulnerability and tool-based control
Robustness and reproducibility of agentic systems
Agentic systems for scientific problems should adapt to different levels of prior knowledge available to the domain (Vinuesa et al., 2024). From the robustness viewpoint, temporal shifts can be counteracted through the use of external resources. Task-oriented specification (Xian et al., 2025) is useful for improving the usability and avoid unnecessary artifacts from model imperfections and the reliability of external tooling and information sources. From the reproducibility viewpoint, agentic tool use should always incorporate essential information of the key parameters. Our work indicates that more research is needed in principled maintenance of agentic frameworks under constant updates of external resources to facilitate reliable agent design (Kapoor et al., 2024).
Limitations
Ethics statement
Appendix A Agentic task
Appendix B Examples of temporal reasoning
(a) No date restriction on tool.
(b) Tool restricted to date before event.
Appendix C Additonal results
Table 3: Performance of ReAct-style LLM agents on text completion ($\gamma=0.75$) following web search with DCTs. For the row of Input, the metrics are computed between the input and the ground truth.
@@ Appendix C Additonal results
We present additional results in Tables 3-4 for the tasks designed for RQ1 and RQ2 described in the main text. These are from experiments carried out at a masking ratio of $\gamma=0.75$. Although most of the performance metrics trend lower than the conditions in Tables 1-2, when $\gamma=0.5$, similar characteristics hold for the time dependence of the STS for publications appeared in more recently (since 2010s) and those before. In the temporal tool selection task, the ReAct + CoT pattern provides performance gain over ReAct-only agents.

### v2

@@ Abstract
Temporal progression is an integral part of knowledge accumulation and update. Web search is frequently adopted as grounding for agent knowledge, yet an improper configuration affects the quality of the agent’s responses. Here, we assess the agent behavior using distinct date-controlled tools (DCTs) as stress test to measure the knowledge variability of large language model (LLM) agents. We demonstrate the temporal effects of an LLM agent as a writing assistant, which uses web search to complete scientific publication abstracts. We show that the temporality of search engine translates into tool-dependent agent performance but can be alleviated with base model choice and explicit reasoning instructions such as chain-of-thought prompting. Our results indicate that agent design and evaluations should take a dynamical view and implement measures to account for the temporal influence of external resources to ensure reliability [^1] The code and datasets for the work are available at https://github.com/RealPolitiX/agent_oost..
@@ 1 Introduction
AI agents based on LLMs and equipped with tools (Mialon et al., 2023; Wang et al., 2024) are well-suited for complex real-world tasks (Gao et al., 2024; Xu et al., 2024) because of their extended capabilities. Their potential to become virtual assistants, paraprofessionals, or “copilots" holds promise for improving the productivity and creativity of the scientific, medical workforces and beyond (Wachter and Brynjolfsson, 2024; Wornow et al., 2024; Bousetouane, 2025). The evaluation standards for AI agents are still in flux (Kapoor et al., 2024; Højmark et al., 2024) and they are urgently needed in specialized domains and realistic scenarios where the the outcomes convey greater bearing on their adoption. Recent works demonstrated the feasibility of LLMs in predicting temporal events (Ye et al., 2024a) and carry out time series forecasting (Tang et al., 2024), but their equivalents in agentic systems are not yet realized. Scientific knowledge and claims have a strong temporal dependence but they have so far been less studied in the context of generative language models (Zhao et al., 2024; Park et al., 2024). We devised a text completion task as a proxy to measure the agent’s usability as a writing assistant with access to external sources (see Fig. 1a).
@@
Web search is an essential tool for grounding agent knowledge to the current and bygone worlds (Pavlick, 2023) and it appears in many applications as a capability extender for models (Zhou et al., 2024a; Song et al., 2024). Nevertheless, web search is subject to the recency and primacy bias of the search engine (Lawrence and Giles, 1998) and the cognitive bias of the users who seek and collect information (Lau and Coiera, 2007). The term search engine manipulation effect (Epstein and Robertson, 2015) was coined to refer to the search results’ influence on public opinions of societal issues. Independent of search engines, factual and scientific knowledge also experience constant but necessary updates over time (Arbesman, 2013).
@@
Figure 1: (a) Illustration of the stress testing framework for agent knowledge, $t_{p}$ indicates the time of publication (b) Temporal tool selection in a ReAct-style agent that performs text completion task in (a) with a selected tool.
@@
While temporal generalization remains challenging for language models (Lazaridou et al., 2021; Wallat et al., 2024), explicit tuning of time-related tool parameters in LLM agents can offer an alternative way to reduce temporal effects of the base model. These effects are a source of performance reliability issues of agentic systems that warrant investigation (Ye et al., 2024b). Date control in reality can manifest passively because the tool-interfaced computer programs have an intrinsic time stamp or a versioned release over time (Zhang et al., 2009). Alternatively, date control can be imposed actively because of copyright, paywall, or local policy. Content access in the past can be controlled retroactively as policy changes (Aral and Dhillon, 2021). From a technical standpoint, invoking different DCTs is equivalent to changing the environment (here means the surface web, see Fig. 1a) of the agent, which requires the agent to adjust to in task execution.
@@
Stress testing is the ultimate test for model behavior and trustworthiness. In the time domain, out-of-sample (OOS) testing is typically used for temporal prediction methods (Hansen and Timmermann, 2015). Analogous OOS assessments in the text domain include predicting future events (Ye et al., 2024a) or generating hypotheses (Zhou et al., 2024b) conditioned on existing (e.g. past) knowledge. We investigate the comparable problem from the tool use perspective, where the agent has access to changing internet-scale information. Because scientific breakthroughs often lead to significant knowledge updates, they are good markers for temporal knowledge progression [^2] Although the judgement on breakthroughs are ultimately subjective and can change with time, our motivation to use them is because of their noticeable footprints on the internet.. In this work, we aim to investigate the following research questions:
@@ 2. RQ2:
    Can agents determine the optimal date-controlled version of a tool to use for a task?
Our contributions along these directions are: (i) Formulation of tool-based stress test for time-dependent knowledge for LLM agents; (ii) Introduction of the SciBreak dataset containing the publication records of public-endorsed scientific breakthroughs from 2000 to 2024. (iii) Investigation of the temporal effects of LLM agent performance and behavior. Besides, we also discuss the impact of temporal information on the agent capability and usability and its implications.
@@ 2 Related works
Previous works have shown that the latent space of LLMs has a direction of time (Gurnee and Tegmark, 2024). Recent investigations show that model performance is affected by the lack of temporal grounding in the pre-training process (Zhao et al., 2024), which can hinder the elicitation of appropriate time-sensitive knowledge at task execution. Previous works have shown LLMs often struggle with tasks that require a consistent temporal grounding (Qiu et al., 2024). The limitation can be improved with techniques such as temporally-informed chain-of-thought (CoT) prompting (Xiong et al., 2024).
@@
3 Tool-based stress testing
@@
Definition 3.3 (Tool-based stress test).
A performance test that induces stress condition by adjusting the tool parameters of an agentic system. A temporal version of the test alters the time information of tools and therefore measures the reliability of agent performance under such conditions.
@@ 4 Testing framework implementation
We constructed the SciBreak dataset, which has a clear time-delimited footprint on the internet—scientific breakthroughs. We extended the dataset collated in Wuestman et al. (2020) to the year of 2024. Each year contains up to $\sim$ 20 publications, including multiple publications contributing to one breakthrough.
Agent configuration
We integrated DCTs into the ReAct (Yao et al., 2023) agent which allows interleaved thinking and action. The agents were constructed from closed-source models, including OpenAI’s GPT-3.5-turbo (gpt-3.5-turbo-0125), GPT-4-turbo (gpt-4-turbo-2024-04-09) (OpenAI, 2024a), and GPT-4o (gpt-4o-2024-08-06) (OpenAI, 2024b) as the base model. For the temporal tool selection task, we also included CoT into the agent pattern (ReAct+CoT) as comparison.
@@
(a) GPT-3.5-turbo
(b) GPT-4-turbo
(c) GPT-4o
Figure 2: Temporal effects of the search engine on agent performance in scientific abstract completion ($\gamma$ = 0.5).
5 Results
Reasoning about time
In our evaluation experiments, the agent’s reasoning behavior related to its awareness of time (e.g. does the tool have time-appropriate utility?) is triggered in two scenarios: (i) When the web search returns nothing or little relevant information to assist task completion. The agent then proceeds to complete the task with the internal knowledge of the base LLM. (ii) In temporal tool selection, when the agent is given an explicit CoT stepwise instruction (Chu et al., 2024) to direct its reasoning towards considering the relavance of information to the topic.
Table 1 show that the STS increases by including CoT prompting (ReAct + CoT) than with ReAct only, when the agent by default selects $\mathcal{T}_{t_{+}-1,t_{+}}$ as the tool. Here, $t\in[t_{-},t_{+}]$ being the date range of the tools. The agent then explores more date choices driven by its internal understanding of the scientific concepts present in the input paragraph. These behavioral characteristics allow the agent to handle non-existent and underspecified contexts in the stress test setting. The performance boost of ReAct + CoT agent pattern requires a model with sufficient reasoning capability such as GPT-4 (OpenAI, 2024a), while for GPT-3.5, it is more prone to failure and the performance gain is reversed.
Table 1: Performance of LLM agents on text completion ($\gamma=0.5$) with temporal tool selection.
Agent model Agent pattern Pub. in 2003 Pub. in 2015
Rouge-L STS Rouge-L STS
GPT-3.5 -turbo ReAct 0.447 0.732 0.447 0.794
ReAct + CoT 0.487 0.640 0.518 0.607
GPT-4 -turbo ReAct 0.635 0.888 0.588 0.783
ReAct + CoT 0.649 0.897 0.644 0.859
Temporal effects across models and masking
High-capacity models with pronounced reasoning capabilities are capable of examining tool dates. The task evaluated the model capability with two different levels of text masking. In Figs. 2-3 and Tables 2-4, the outcome contains two major trends: (i) The more advanced models can recover more of the missing semantic content in the masked input, as indicated by the significant increase of STS from LLM agents based on GPT-3.5 to GPT-4o (OpenAI, 2024b). (ii) There is noticeable variability of agent performance between knowledge generated more recently than before 2010. (iii) For the same model, varying the masking ratio of input largely preserves the date sensitivity in the performance. Similar time-dependent performance change has been described in a different context for LLMs (Zhao et al., 2024). Overall, the temporal effects are less severe in more capable models.
6 Discussion
Agent vulnerability and tool-based control
Robustness and reproducibility of agentic systems
Agentic systems for scientific problems should adapt to different levels of prior knowledge available to the domain (Vinuesa et al., 2024). From the robustness viewpoint, temporal shifts can be counteracted through the use of external resources. Task-oriented requirements specification (Xian et al., 2025) is useful for improving the usability and avoid unnecessary artifacts from model imperfections and the reliability of external tooling and information sources. From the reproducibility viewpoint, agentic tool use should always incorporate essential information of the key parameters. Our work indicates that more research is needed in principled maintenance of agentic frameworks under constant updates of external resources to facilitate reliable agent design (Kapoor et al., 2024).
Limitations
Ethics statement
Appendix A Agentic task
Appendix B Dataset preprocessing
The SciBreak dataset is partly based on peer-reviewed publications collated and categorized in Wuestman et al. (2020), including the annual top-ten-ranked scientific breakthroughs from mid-1990s till 2012 collated by the journal Science at the end of each year. The publications are drawn from various journals in the physical, biomedical, and engineering sciences, which constitute the scope of the ranking. We chose records from year 2000 to 2012 and extended to year 2024 by self-curating the extra years of ranked publications from the published tally in each year. The links to the yearly breakdown is provided as follows: 2013, 2014, 2015 (12), 2018 (17), 2021 (14), 2024 (11).
We collected the abstracts of the associated publications through web scraping from the public databases PubMed [^3] https://pubmed.ncbi.nlm.nih.gov/ and SAO/NASA ADS Abstract Service [^4] https://ui.adsabs.harvard.edu/ using the Digital Object Identifiers of the publications, which are also provided in the dataset.
Appendix C Extended results
Table 2: Performance of ReAct-style LLM agents on text completion ($\gamma=0.5$, see Fig. 2) following web search with DCTs. For the row of Input, the metrics are computed between the input and the ground truth.
@@ Appendix C Extended results
Table 3: Performance of ReAct-style LLM agents on text completion ($\gamma=0.75$, see Fig. 3) following web search with DCTs. For the row of Input, the metrics are computed between the input and the ground truth.
@@
(a) GPT-3.5-turbo
(b) GPT-4-turbo
(c) GPT-4o
Figure 3: Temporal effects of the search engine on agent performance in scientific abstract completion ($\gamma$ = 0.75).
Extended results for RQ1 include Table 2, which shows a portion of the results in Fig. 2, and Table 3, which is similarly related to Fig. 3. Results in Table 4 are the extension for temporal tool selection (RQ2) with a different masking ratio (compare Table 1). These are from experiments carried out at a masking ratio of $\gamma=0.75$. Although most of the performance metrics trend lower than the conditions in Tables 2-1, when $\gamma=0.5$, similar characteristics hold for the time dependence of the STS for publications appeared in more recently (since 2010s) and those before. In the temporal tool selection task, the ReAct + CoT pattern provides performance gain over ReAct-only agents.
(a) No date restriction on tool.
(b) Tool restricted to date before event.
Appendix D Examples of temporal awareness